# BenchPress: A Human-in-the-Loop Annotation System for Rapid Text-to-SQL Benchmark Curation

### v1

@@ Abstract.
To address this challenge, we introduce BenchPress, a human-in-the-loop system designed to accelerate the creation of domain-specific text-to-SQL benchmarks. Given a SQL query, BenchPress uses retrieval-augmented generation (RAG) and LLMs to propose multiple natural language descriptions. Human experts then select, rank, or edit these drafts to ensure accuracy and domain alignment. We evaluated BenchPress on annotated enterprise SQL logs, demonstrating that LLM-assisted annotation drastically reduces the time and effort required to create high-quality benchmarks. Our results show that combining human verification with LLM-generated suggestions enhances annotation accuracy, benchmark reliability, and model evaluation robustness. By streamlining the creation of custom benchmarks, BenchPress offers researchers and others a mechanism for assessing text-to-SQL models on a given domain-specific workload. BenchPress is freely available via our public GitHub repository [^1] https://github.com/fabian-wenz/enterprise-txt2sql and accessible for use on our website [^2] http://dsg-mcgraw.csail.mit.edu:5000/.
@@ 1. Introduction
To address this gap, we introduce BenchPress, a system designed to enable organizations to create their own text-to-SQL benchmarks quickly and efficiently. By combining LLM-generated suggestions with human validation, BenchPress produces accurate, domain-specific training data while reducing annotation effort. It offers a scalable and privacy-aware foundation for enterprise-grade text-to-SQL evaluation.
BenchPress operates in one primary way:
SQL-to-NL generation
Given a SQL query, BenchPress generates a natural language (NL) description, allowing domain experts to review and refine it, significantly reducing annotation effort.
@@ 2. Related Work
Unlike prior systems, BenchPress provides a practical toolkit for constructing domain-specific text-to-SQL benchmarks. It supports now mainly SQL-to-NL annotation for validation, semantic enrichment, and human verification, enabling scalable, robust benchmarking in enterprise settings. BenchPress provides a means to generate new, customized corpora—tailored to real-world workloads.
@@ 3. Enterprise SQL Logs: Challenges
In this research, we have worked with four text-to-SQL benchmarks: Spider, Bird, and Fiben as public datasets, and Beaver as a private, enterprise-oriented dataset. Beaver is based on SQL logs from enterprise databases across industries such as education, technology, and manufacturing. These logs span in total over 300 schemas and nearly 4000 queries, featuring complex, multi-source schemas with inconsistent naming conventions and semantic overlaps. For example, in the MIT data warehouse, a single natural language query can often be answered by multiple SQL queries due to the presence of materialized views and semantic ambiguity across tables.
Next we discuss the challenges of working with enterprise SQL logs in contrast to those from open-domain benchmarks:
@@
Queries in enterprise settings are substantially more complex than those in public benchmarks. They commonly include nested subqueries, aggregation dependencies, and recursive joins. A single query may aggregate data from 5–10 tables and use the result in a nested filter. LLMs, which are typically trained on simpler academic benchmarks, struggle to handle this structural depth. A comparative graph illustrating query complexity follows.
@@
Enterprise data often contains sensitive or proprietary information, such as employee salaries or internal metrics. Due to strict privacy constraints, this data cannot be publicly released or used for model training. BenchPress addresses this challenge by enabling organizations to evaluate public models securely on private SQL logs—supporting informed model selection without compromising confidentiality.
@@
Enterprise data warehouses often aggregate tables from various internal systems, leading to schema inconsistencies. It is common to find multiple tables with identically named columns like “user_id” that actually refer to different entities. Disambiguating such cases requires either careful schema documentation or intelligent annotation, both of which pose a challenge for LLM-based systems.
@@
Many enterprise datasets suffer from sparsity and imbalance. For example, the Intel data warehouse includes performance metrics from millions of devices, but many fields may be missing or unevenly distributed. LLMs trained on uniformly structured or synthetic data may struggle with these irregularities, resulting in biased performance on common patterns and failures on rare or incomplete cases.
@@ 4.1. Workflow Overview
    Project Setup: The user selects or creates a new annotation project tied to a specific enterprise workload.
@@
    Data Ingestion: Users upload SQL logs and schema files or select from one of four supported public benchmarks: Bird, Fiben, Spider, or Beaver. The system parses and stores this data for further processing.
@@
    Task Configuration: Users choose the annotation direction (currently only SQL-to-NL), as well as the language model (e.g., GPT-4o, GPT-3.5 Turbo, or DeepSeek).
Annotation Loop:
@@
    (Optional) Decomposition: For nested SQL queries, the system rewrites the query into a series of Common Table Expressions (CTEs), breaking it down into logically independent subqueries.
@@
    Feedback: Annotators can rank, refine, or discard priorities assigned to the LLM. This human-in-the-loop feedback improves prompt quality over time and supports future fine-tuning Christiano et al. (2017).
@@ 4.2. Key Design Features
  • Human-in-the-loop Feedback: At its core, BenchPress emphasizes continuous human oversight. Domain experts iteratively review, edit, rank, and flag unclear annotations produced by the model. This structured, iterative review phase ensures annotations meet enterprise-quality standards and reduces error propagation downstream. This reflects Google’s "PAIR" principles Dixon and Terry (2023) for responsible AI and supports findings that combining human judgment with AI outputs results in higher quality, robustness, and trustworthiness Amershi et al. (2019).
@@ 5.1. Setup
To evaluate the impact of BenchPress on annotation efficiency and quality, we conducted a controlled user study using a between-subjects experimental design. In this design, each participant is randomly assigned to one condition only—ensuring that comparisons across conditions reflect differences in the interface or workflow, rather than learning effects or fatigue. This setup is widely used in HCI and behavioral research, including in experimental frameworks Lazar et al. (2017).
A total of 18 participants were recruited and first grouped into two strata—advanced and non-advanced SQL users—based on a pre-study questionnaire assessing their experience and familiarity with relational databases. Within each stratum, participants were randomly assigned to one of three experimental conditions using a balanced Latin square design to ensure counterbalancing:
@@
Each participant was assigned the same set of 30 SQL queries, sampled from the Beaver and Bird datasets (anonymized), and instructed to write a natural language description for each SQL query. The task was SQL-to-NL annotation only.
Independent variables in the study were the annotation condition (BenchPress, Manual, LLM) and user expertise (Advanced, Non-Advanced). The primary dependent variables included annotation time and annotation quality. Quality was assessed using both observational measures (e.g., back-translation match, ROUGE similarity) and participant self-reports on task confidence and perceived difficulty.
This setup enables a structured comparison of how different interfaces and user backgrounds affect performance in enterprise SQL annotation tasks.
@@ 5.2. Evaluation
We evaluated the performance of each annotation condition—BenchPress, Manual, and LLM—across three dimensions: annotation accuracy, annotation latency, and semantic fidelity using a backtranslation task.

### v2

@@ Abstract.
To address this challenge, we introduce BenchPress, a human-in-the-loop system designed to accelerate the creation of domain-specific text-to-SQL benchmarks. Given a SQL query, BenchPress uses retrieval-augmented generation (RAG) and LLMs to propose multiple natural language descriptions. Human experts then select, rank, or edit these drafts to ensure accuracy and domain alignment. We evaluated BenchPress on annotated enterprise SQL logs, demonstrating that LLM-assisted annotation drastically reduces the time and effort required to create high-quality benchmarks. Our results show that combining human verification with LLM-generated suggestions enhances annotation accuracy, benchmark reliability, and model evaluation robustness. By streamlining the creation of custom benchmarks, BenchPress offers researchers and others a mechanism for assessing text-to-SQL models on a given domain-specific workload, where no comparable tools exist today. BenchPress is freely available via our public GitHub repository [^1] https://github.com/fabian-wenz/enterprise-txt2sql and accessible for use on our website [^2] http://dsg-mcgraw.csail.mit.edu:5000/.
@@ 1. Introduction
To address this gap, we introduce BenchPress, a novel system designed to enable organizations to create their own text-to-SQL benchmarks quickly and efficiently. Unlike prior work, BenchPress focuses on the SQL-to-NL annotation problem itself—a capability largely absent in existing tools and particularly important in enterprise settings, where SQL logs are abundant but corresponding natural language descriptions are scarce. Given a SQL query, BenchPress generates a natural language (NL) description, allowing domain experts to review and refine it, significantly reducing annotation effort. It offers a scalable and privacy-aware foundation for enterprise-grade text-to-SQL evaluation.
@@ 2. Related Work
Unlike prior systems, BenchPress provides a practical toolkit for constructing domain-specific text-to-SQL benchmarks. The current version primarily focuses on SQL-to-NL annotation, which serves multiple downstream purposes including validation, semantic enrichment, and human verification; thereby enabling scalable and robust benchmarking in enterprise settings. BenchPress provides a means to generate new, customized corpora tailored to real-world workloads.
@@ 3. Enterprise SQL Logs: Challenges
Query Sets #Keywords #Tokens #Tables #Columns #Agg #Nestings
BEAVER (DW) 15.6 99.8 4.2 11.9 5.5 2.05
Spider $\downarrow$ 80.8% $\downarrow$ 81.5% $\downarrow$ 64.3% $\downarrow$ 75.6% $\downarrow$ 83.6% $\downarrow$ 45.5%
FIBEN $\downarrow$ 39.1% $\uparrow$ 62.2% $\downarrow$ 9.5% $\downarrow$ 18.5% $\downarrow$ 63.6% $\downarrow$ 23.8%
BIRD $\downarrow$ 73.1% $\downarrow$ 68.7% $\downarrow$ 54.7% $\downarrow$ 63.0% $\downarrow$ 87.3% $\downarrow$ 45.5%
Table 1. Query-level complexity metrics across benchmarks:
This table compares structural and linguistic query characteristics between BEAVER’s Datawarehouse (DW) and existing benchmarks (Spider, FIBEN, BIRD). Percentages indicate the relative difference compared to BEAVER (DW).
Data Sets Columns/Table Rows/Table Table/DB Uniqueness Sparsity Data Types
BEAVER (DW) 15.6 128K 99 45.9% 15.0% 4
Spider $\downarrow$ 65.4% $\downarrow$ 98.4% $\downarrow$ 94.8% $\uparrow$ 59.5% $\downarrow$ 100% 0.0%
FIBEN $\downarrow$ 84.0% $\downarrow$ 40.6% $\uparrow$ 53.5% $\uparrow$ 28.1% $\downarrow$ 100% $\uparrow$ 100%
BIRD $\downarrow$ 56.4% $\uparrow$ 328.9% $\downarrow$ 54.7% $\uparrow$ 72.8% $\downarrow$ 100% $\uparrow$ 74%
Table 2. Data-level complexity metrics across benchmarks:
This table reports schema and data characteristics comparing BEAVER’s DW to Spider, FIBEN, and BIRD. Percentages show the relative difference compared to the DW section of BEAVER.
In this research, we have worked with four text-to-SQL benchmarks: Spider, Bird, and Fiben as public datasets, and Beaver as a private, enterprise-oriented dataset. Beaver is based on SQL logs from enterprise databases across industries such as education, technology, and manufacturing. These logs span in total over 300 schemas and nearly 4000 queries, featuring complex, multi-source schemas with inconsistent naming conventions and semantic overlaps. For example, in the MIT data warehouse, a single natural language query can often be answered by multiple SQL queries due to the presence of materialized views and semantic ambiguity across tables. This is visualized in Tables 2 and 2.
In Tables 2 and 2, the arrows denote the relative change of each benchmark with respect to the Beaver (DW) workload: a downward arrow ($\downarrow$) indicates a decrease and an upward arrow ($\uparrow$) indicates an increase relative to Beaver (DW). The color coding highlights whether this change represents an easier (red) or harder (green) characteristic for the text-to-SQL task. Importantly, higher values are not universally “better” or “worse”; their desirability depends on how they influence model difficulty.
Next, we discuss the challenges of working with these enterprise SQL logs in contrast to those from open-domain benchmarks:
@@
Queries in enterprise settings are substantially more complex than those found in public benchmarks. They commonly include nested subqueries, aggregation dependencies, and recursive joins. A single enterprise query may aggregate data from 5–10 tables and use the result in a nested filter. LLMs, which are typically trained on simpler academic benchmarks, struggle to handle this structural depth. In Table 2, higher counts of keywords, nestings, aggregations, and referenced tables all quantify this complexity. Each of these dimensions reflects a form of compositional depth that increases the cognitive load for annotators and the reasoning burden on LLMs, contributing to the substantial enterprise gap observed across workloads.
@@
Enterprise data often contains sensitive or proprietary information, such as employee salaries or internal performance metrics. Because of strict privacy constraints, this data cannot be and traditionally has not been publicly released, which in turn makes it challenging to build models, create representative benchmarks, and evaluate performance accurately. Consequently, organizations must evaluate any model in a secure, private environment using their own in-domain data before deployment.
@@
Enterprise data warehouses often aggregate tables from various internal systems, leading to schema inconsistencies. It is common to find multiple tables with identically named columns such as “user_id” that actually refer to different entities. Disambiguating such cases requires either careful schema documentation or intelligent annotation, both of which pose challenges for LLM-based systems. In Table 2, lower schema uniqueness, i.e., more repeated or semantically similar column names across tables, directly reflects this ambiguity. Because models must distinguish between near-identical attributes, a decrease in uniqueness ($\downarrow$) is marked as green, indicating increased benchmark difficulty. Similarly, lower data-type diversity amplifies ambiguity by reducing the semantic cues available for distinguishing column roles; when tables use only few data types (e.g., diversity near 0.0), LLMs lose an important signal for disambiguation.
@@
Many enterprise datasets suffer from sparsity and imbalance. We observe this not only in Beaver, but also in a large-scale Intel-internal data warehouse, in which real-world operational data exhibits uneven field completeness, heterogeneous logging practices, and highly skewed attribute distributions. For example, the Intel data warehouseZhu et al. (2025) includes performance metrics from millions of devices, yet many fields may be partially or entirely missing. LLMs trained on uniformly structured or synthetic data often struggle with these irregularities, resulting in biased performance on common patterns and failures on rare or incomplete cases. In Table 2, these effects are captured by higher schema sparsity, meaning a greater proportion of NULL or missing values per column. Additionally, a higher number of columns per table and a larger number of rows per table indicate greater structural and statistical imbalance: models must process wider and deeper tables, often with uneven population across attributes. All of these factors increase annotation and modeling difficulty and are therefore marked as preferable ($\uparrow$) in green for benchmarking purposes.
@@ 4.1. Workflow Overview
    Project Setup: The API key is kept entirely within the user’s browser storage, ensuring that it never leaves the local client. The username serves as a local workspace identifier under which annotation projects are organized. Users can maintain multiple projects, each associated with a specific schema and the SQL logs or queries they upload.
@@
    Dataset Ingestion: Users upload SQL logs and schema files or select from one of four supported benchmarks: Bird, Fiben, Spider, or Beaver. Because SQL logs and schemas can be large, they are stored on the server rather than in the browser. Standard browser storage mechanisms such as localStorage typically allow only about 5–10 MB per origin  mdn (b), and although IndexedDB can support larger datasets, the quota is highly variable across browsers and cannot guarantee availability for large enterprise SQL logs  mdn (a). Centralized server-side storage ensures the full dataset is always available for retrieval-augmented generation (RAG), which requires global access to all uploaded documents for vector search and efficient retrieval. The system then parses and stores this data for further processing.
@@
    Task Configuration: Users choose the annotation direction task (currently only SQL-to-NL), as well as the language model (e.g., GPT-4o, GPT-3.5 Turbo, or DeepSeek).
Annotation Loop: To illustrate how a single query progresses through the annotation loop, Figure 3 shows a concrete example passing through decomposition, retrieval, candidate generation, recomposition, and feedback.
Figure 3. Example annotation run in BenchPress. This figure illustrates how a single SQL query moves through the annotation loop. For nested queries, BenchPress performs optional decomposition into CTEs, followed by schema and example retrieval, LLM-based NL candidate generation, optional recomposition into a unified explanation, and human feedback. This example corresponds directly to steps 3.5–6 in the Annotation Loop.
@@
    (Optional) Decomposition: For nested SQL queries, the system rewrites the query into a series of Common Table Expressions (CTEs), breaking it down into semantically logical subqueries.
@@
    Feedback: Annotators can rank, refine, discard, or add priorities assigned to the LLM, and can also introduce external domain knowledge or highlight common failure patterns that the model typically misses. This human-in-the-loop feedback loop not only corrects the LLM’s current reasoning but also enriches it with enterprise-specific insights, thereby improving prompt quality over time and providing strong supervision for future fine-tuning Christiano et al. (2017).
@@ 4.2. Key Design Features
  • Human-in-the-loop Feedback: At its core, BenchPress emphasizes continuous human oversight. Domain experts iteratively review, edit, rank, discard, or add priorities to the annotations proposed by the LLM, and can inject external domain knowledge or highlight common failure modes that the model typically misses. This structured, iterative feedback loop ensures that annotations meet enterprise-quality standards and prevents error propagation downstream. The approach aligns with Google’s PAIR principles for responsible AI Dixon and Terry (2023) and is consistent with findings that combining human judgment with AI outputs leads to higher quality, robustness, and trustworthiness Amershi et al. (2019).
@@ 5.1. Setup
To evaluate the impact of BenchPress on annotation efficiency and quality, we conducted a controlled user study using a between-subjects experimental design, meaning that each participant experiences only one of the available conditions. This design prevents learning effects, familiarity with the queries, or fatigue from carrying over between interfaces. In this design, each participant is randomly assigned to one condition only, ensuring that observed performance differences reflect the properties of the interface rather than improvements from repeated exposure. This setup is widely used in HCI and behavioral research, including in experimental frameworks Lazar et al. (2017).
Our study includes 18 participants, which is well within the range commonly accepted in HCI and empirical systems research. Prior work in usability testing has repeatedly shown that relatively small samples (5–20 participants) are sufficient to uncover the vast majority of issues and yield stable performance differences between interfaces. Nielsen’s foundational work demonstrates that diminishing returns set in rapidly after 15 participants Nielsen (1994), while Virzi provides empirical evidence that additional users contribute only marginal increases in explanatory power Virzi (1992). Faulkner further shows that studies with 5–20 users achieve stable, reliable results across usability metrics Faulkner (2003). Similarly, the “small-N, high-signal” evaluation paradigm used throughout systems-HCI research (e.g., in CHI, UIST, and CSCW) relies on focused studies with 12–20 participants to draw statistically meaningful conclusions about task performance, efficiency, and error rates.
These 18 participants were then first grouped into two strata – advanced and non-advanced SQL users-based on a pre-study questionnaire assessing their experience and familiarity with relational databases. Within each stratum, participants were randomly assigned to one of three experimental conditions using a balanced Latin square design to ensure counterbalancing:
@@
Each participant was assigned the same set of 30 SQL queries, sampled from the Beaver and Bird datasets, and instructed to write a natural language description for each SQL query. The task was SQL-to-NL annotation only. A practical constraint in this setting is the cold-start condition: at the beginning of the study, no prior annotations exist, and thus the system cannot retrieve similar examples via RAG. Consequently, the first annotation is performed without example-based guidance, while subsequent annotations incrementally benefit from the examples created earlier in the session and the knowledge injected in the feedback-loop. This mirrors realistic enterprise deployments, where organizations typically begin with no internal annotation archive, and manually building such an example set would itself require substantial effort. BenchPress therefore assumes a cold start by default, while in future enterprise use cases we intend to offer both options: starting from an empty knowledge base or initializing the system with examples from public text-to-SQL datasets.
Independent variables in the study were the annotation condition (BenchPress, Manual, Vanilla LLM) and user expertise (Advanced, Non-Advanced). The primary dependent variables included annotation time and annotation quality. Quality was assessed using observational measures (e.g., back-translation match, ROUGE similarity). This setup enables a structured comparison of how different interfaces and user backgrounds affect performance in enterprise SQL annotation tasks.
@@ 5.2. Evaluation
We evaluated the performance of each annotation condition – BenchPress, Manual, and Vanilla LLM – across three dimensions: annotation accuracy, annotation latency, and semantic fidelity using a backtranslation task.
@@ 5.3. Summary and Observations
6. Enterprise SQL Logs: Contributions
In this section, we highlight the contributions how BenchPress addresses challenges unique to enterprise SQL logs explicit by discussing (i) schema complexity, (ii) domain-specific terminology, and (iii) privacy and confidentiality constraints. As introduced in Section 1, these challenges are well-documented in enterprise workloads such as Beaver, and they directly inform the design of BenchPress.
Schema Complexity and Ambiguity
As discussed in Section 3, enterprise schemas often exhibit extensive duplication of names, values, and attribute types. This kind of structural redundancy and overlap is characteristic of large warehouse systems and rarely appears in public academic datasets. BenchPress mitigates this challenge by retrieving past SQL queries that use the same columns, thereby exposing annotators to the actual join patterns and filters associated with ambiguous attributes. For example, retrieved usage clarifies that ACADEMIC_TERMS_ALL contains both historical and current term records, whereas ACADEMIC_TERMS focuses only on the current period. Similarly, filters such as STREET_TYPE = ’STREET’ in address tables indicate that only physical street addresses—not mailing addresses—should be considered. By surfacing these real usage patterns, BenchPress turns otherwise opaque schemas into interpretable annotation context and reduces confusion that would otherwise propagate into downstream benchmarks.
Domain-Specific Terminology and Abbreviations
Enterprise SQL logs also contain specialized vocabulary from domains such as finance, HR, and student administration—terminology that is often absent from table documentation. BenchPress assists annotators by generating multiple candidate NL descriptions informed by retrieved query contexts. These candidates help annotators quickly infer the correct domain meaning while avoiding hallucinations, enabling effective annotation even when no glossary or documentation is available—a condition frequently noted by reviewers.
Privacy and Confidentiality Constraints
Enterprise SQL logs also contain specialized vocabulary from domains such as finance, HR, and student administration—terminology that is often absent from table documentation. BenchPress assists annotators by generating multiple candidate NL descriptions informed by retrieved query contexts. These candidates help annotators quickly infer the correct domain meaning while avoiding hallucinations, enabling effective annotation even when no glossary or documentation is available—a condition frequently noted by reviewers. Moreover, through BenchPress’s iterative feedback loop, any domain-specific clarifications provided by annotators are fed back into the prompt for subsequent queries. This eliminates the need to repeatedly look up the same information in external documentation, as once a piece of knowledge has been captured, the system is able to reuse it automatically in future generations.
Connection to Enterprise Difficulty
The Beaver benchmark demonstrates that state-of-the-art LLMs suffer a substantial execution accuracy drop (up to 70–90%) when evaluated on enterprise SQL logs, despite strong performance on public datasets. Tables 2 and 2 quantify these differences in query and data complexity.
The Beaver benchmark shows that state-of-the-art LLMs experience dramatic execution accuracy drops on enterprise workloads due to schema complexity, domain terminology, and data irregularity. Tables 2 and 2 illustrate these differences. BenchPress is designed explicitly to help organizations create domain-specific text-to-SQL benchmarks that reflect these real-world complexities, thereby reducing deployment risks and improving model evaluation fidelity.
@@ 7. Conclusions and Future Work
Acknowledgments
We thank Michael Stonebraker, Peter Baile Chen, Bijan Arbab, Moh Haghighat and Layne Mills for their invaluable feedback throughout the development of BenchPress. Although not listed as co-authors due to potential conflicts with other CIDR submissions, their contributions were instrumental to this work. We also thank the participants of our user study for their time and effort, which helped us evaluate and refine the system. This research is supported by Intel as part of the MIT Data Systems and AI Lab (DSAIL).